Status Report

NASA Space Station On-Orbit Status 11 Jan 2004

By SpaceRef Editor
January 11, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 11 Jan 2004

All ISS systems continue to function nominally, except as noted previously or below.  Ahead: Week 12 for Increment 8.

Sunday — but not quite the usual rest day for the crew (neither was Saturday, yesterday).

Leak Test Activities — Update and Preview (ref. On-Orbit Status report 1/10):
The crew eliminated Progress 12P and its docking assembly yesterday as a source of the pressure decay.   Today, isolation/monitoring tests are focusing on the U.S. Airlock (A/L) and the combined DC-1/Soyuz volume.  Depending on the results, which should be available by tonight, ground specialists & crew will either have found the source or press on to prepare for isolation of the RS and USOS segments next Wednesday, as planned (final details are still being worked).  Regardless of today’s checks and their outcome, ISS will be repressed tonight, just prior to crew sleep, using Progress O2 (not N2 from A/L HPGTs).   [At 3:40am EST this morning, ppN2 was 558.3 mmHg (torr), a drop of ~1.12 mmHg (= 0,022 psi) over 24 hours.  This is a decay of 0.2% from yesterday’s ppN2, or a total loss of 3% from the nominal ppN2 of ~576 mmHg.]

Early in the morning, CDR/SO Michael Foale began with preparations for the isolation testing of the A/L “Quest”.  Atmospheric pressure inside the isolated A/L is now being monitored by the ground for several hours.   [The work involved turning on the four A/L GLAs (general luminaire assembly) lighting fixtures, closing the depress pump valve, installing a VRA (vent relief valve) power cable, removing the PCS (portable computer system) laptop, other equipment and personal items, and closing the Node starboard hatch.]

FE Alexander Kaleri, in the RS, similarly prepared the Soyuz/DC-1 “Pirs” combination for the leak monitoring, which he is conducting throughout the day, once every hour for a duration of ten hours.  [The two hatches between DC-1 and Soyuz TMA-3 are being kept open to retain quick access safety.  Tonight at ~4:00pm EST, Sasha will restore the initial configuration.]

Kaleri completed the weekly routine life support systems (SOZh) maintenance tasks of inspection and collecting SP toilet flush counter and SVO water supply readings in the Service Module (SM) for calldown to TsUP.  He also was scheduled to do the regular inspection of the Elektron oxygen generator’s VM gas/liquid system (GZhS) for obstructing air bubbles. 

There were no SFOG (solid-fuel oxygen generator) candles burnt today, so as not to disturb the ongoing pressure monitoring.  The number of candles used so far is 20 (2 units/day), with 122 SFOGs (Russian: TGK) remaining on board.  [New SFOGs have been built by Russia and are currently in safety certification.]

Mike concluded the current renal (kidney) stone experiment session by collecting and stowing equipment and samples of the last two days.   [The ground expressed a heartfelt Thank you for the crew’s continued support for this experiment.  There will be one more in-flight session for Mike and Sasha in April, and they were reminded to keep taking their pills.]

The crew worked out on TVIS, RED exerciser and VELO cycle with load trainer.

At 4:25 pm, just prior to crew sleep, on TsUP “Go”, ISS will be repressed using Progress O2 (not N2 from U.S. A/L supplies). [The plan is to open the Progress O2 valve and leave it open through the night.  There is approximately 8-10 mmHg of O2 remaining in the Progress and this repress will bring the cabin safely above the 13.9 psi certification limit on some hardware.]

Upcoming Events:

  • 12P Undock — 1/28/04
  • 13P Launch — 1/29/04
  • 13P Dock — 1/31/04

Weekly Science Update (Expedition Eight — 10th):
GASMAP:   Thank you to the crew for supporting another 30-day health check, with flawless execution, in association with collecting environmental sampling data for the US Lab.  The data collected has been forwarded to ECLSS for processing and analysis.

Human Research Facility/Workstation (HRF WS):  Continuing.

Advanced Ultrasound:  Planned.

Hand Posture Analyzer (HPA): Mike’s second (of three) HPA session is scheduled for next week.  The ground has analyzed the video of the first session and this has helped to evaluate acquired data correlated with pictures. 

In-Space Soldering Investigation (ISSE): Thanks to Mike for wrapping the ISSI coupons in preparation for the soldering operations later in the Increment.  The ground enjoyed watching him and appreciates his dedication to details.

Pore Formation and Mobility Investigation (PFMI):   The PFMI experiment conducted on1/8 went great.  Both the sample and camera translation systems operated without any problems and even the video was as good as any that has been seen on the ground.  The sample processed was an alloy sample, and investigators were able to observe transitions from planar to cellular to dendritic growth. The time to change from a planar to a cellular growth front was observed to take much longer in micro-G than in a sample processed under essentially identical conditions on earth. PFMI is back in business.
Space Acceleration Measurement System (SAMS):   The SAMS file repair procedure that was performed on 1/5 was successful, and SAMS was able to collect data all week.  SAMS successfully supported the PFMI operations this week.  Upcoming events include additional support for PFMI.
Microgravity Acceleration Measurement System (MAMS):  MAMS continues to measure the quasi-steady and vibratory environment of the ISS.  Reboost and recent attitude changes were captured and are being analyzed.    

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Behaving nominally.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):   Planned.

Renal Stone (RS): Thanks to the crew for taking the time out of their hectic week to provide an accurate hardware inventory for the Renal Stone experiment.  The ground is looking forward to hearing how the second inflight collections ended.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):  Pre-Increment requirements have been completed.  A second BBT (Beacon & Beacon Tester) session will be scheduled in US Lab.

Foot/Ground Reaction Forces during Space Flight (FOOT):  The ground looks forward to another excellent data collection.

Materials ISS Experiment (MISSE):  In progress. Deployed outside. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI): The CBOSS-FDI team and the NIH (National Institute of Health) PI sincerely appreciate all of the crew’s efforts on FDI this week.  They are still in the process of analyzing the video and images that the crew took throughout last week.  The new setup is considered a success and it is hoped that it makes procedures a bit easier and faster to manage.  The first images were just a bit too bright, which might have been due to the new MWA (maintainance work area) setup; the lower f-stop should help out in the analysis.  Investigators look forward to the data coming down from the 1/9 activities.

Earth Knowledge Acquired by Middle School Students (EarthKAM): Students are looking forward to the next session later this Increment.

Earth Science Toward Exploration Research (ESTER):  Looking forward to the sessions next year.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2): Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):   Completed.

Miscible Fluids in Microgravity (MFMG):    Mike was thanked for setting up MFMG this week.  The team is looking forward to MFMG thermal pperations using the CGBA the week of 1/20.

Educational Payload Operations (EPO):  The Educators are excited about using the video from the crew’s education demonstrations in classrooms and workshops.  They were thanked for providing this valuable education resource.  The ground is looking forward to upcoming activities.

Crew Earth Observations (CEO):  The ground now has a significant backlog of imagery to review and catalog.  It was noticed that the crew seem to have been experimenting with varying focal lengths during some of their sessions at the window.  This can be helpful in locating and composing a good image of a particular target.  However, they should remember that there are usually specific lens recommendations for most of the targets and that changing focal length usually requires refocusing.  The crew was invited keep this aspect of the camera system in mind.  They now seem to have become so familiar with both the selected targets and the features/phenomena they represent, that they are now acquiring them or their proxies even when not solicited.

Today’s optional CEO targets, in the current XPOP attitude constrained by flight rule to fewer near-vertical targets due to shutter closure, were S Chad swamplands(sunglint opportunity left of track.  Early December CEO images of this area show many changes compared with the Landsat [1990] images.  Sunglint images will reveal which streams in the complex network are presently active. Sediment dispersal patterns [i.e. river patterns] on continents are a new topic of geologic interest.  Large modern continental sediment bodies like those of the swamplands had never been examined systematically by geologists until handheld imagery provided the concept for a global study),Congo-Ubangui R. confluence (Dynamic event.  Unusual clearing in the northern Congo basin: Congo River water is dark and tannin stained forest-derived water without significant sediment; the Ubangui R. is light colored with sediment from the semiarid plains of the Sahel.  Looking right for the confluence), and Panama (Dynamic event. Nadir pass. Clear skies may have persisted for a view of this seldom seen point).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

SpaceRef staff editor.